OpenStudy (anonymous):

what is the anti derivative of 4/(1+4t^2)

OpenStudy (australopithecus):

you can use trigonometric substitution for this

OpenStudy (anonymous):

didnt recognize that , can you show me?

OpenStudy (australopithecus):

this lists the three cases https://en.wikipedia.org/wiki/Trigonometric_substitution In this case you would use tan(x)

OpenStudy (anonymous):

great more formulas to remember before the test wednesday :)

OpenStudy (anonymous):

could i also use u substitution for this?

OpenStudy (australopithecus):

so first factor out 4 4/4(1/4 + t^(2)) = 1/(1/4 + t^(2)) now using the identity tan^(2)(theta) + 1 = sec^2(theta) so we have that t = (1/4)^(1/2)tan(theta) t = (1/2)tan(theta) Now solve for theta theta = arctan(2t) Now you need to right the integral it in terms of theta so d(theta) = (1/2)sec^(2)(theta)dx

OpenStudy (australopithecus):

no you cannot use u substitution for this

OpenStudy (anonymous):

oh god we havent done much trig substiution if any..

OpenStudy (australopithecus):

that 1/4 has to be 1 + tan^(2)(theta) crap

OpenStudy (anonymous):

is there any other way to do this problem?

OpenStudy (australopithecus):

lol this only looks hard it really isn't and no this is a special case

OpenStudy (australopithecus):

I guess you could use partial fractions

OpenStudy (australopithecus):

Really you should just watch a video on trig sub, try khan academy he explains it really well

OpenStudy (anonymous):

thanks will do

OpenStudy (australopithecus):

u = 1+4t^2 du = 8tdx du/8t = dx \[\int\limits_{}^{} 4du/(u)8t\] yeah this will get ugly if we use u sub as (u - 1)/4 =t^(2) ((u-1)/4)^(1/2) = t

OpenStudy (australopithecus):

I messed up on my algebra and my brain is too dead I wil lget someone else to help you

OpenStudy (australopithecus):

This video might apply to oyur problem http://www.khanacademy.org/math/calculus/v/integrals--trig-substitution-2

OpenStudy (australopithecus):

I think I made a mistake factoring the problem initially let me try again

OpenStudy (anonymous):

ok

OpenStudy (australopithecus):

so \[\int\limits_{}^{}4dt/(1 + 4t^{2})\] so set 2t = tan(theta) t = tan(theta)/2 theta = arctan(2t) Now write it in terms of d(theta) d(theta) = sec^(2)(theta)/2 dt 2d(theta)/sec^(2)(theta) = dt \[\int\limits_{}^{}d(\theta)8/\sec^{2}(\theta)(1+\tan^{2}(\theta)) = 8\int\limits_{}^{}d(\theta)/\sec^{4}(\theta)\]

OpenStudy (australopithecus):

\[8\int\limits_{}^{}d(\theta)/(1/\cos(\theta)^{4}) = 8\int\limits_{}^{}\cos^{4}(\theta)d(\theta)\]

OpenStudy (australopithecus):

this can be solved pretty easily using identities now

OpenStudy (anonymous):

thanks think i got it now

OpenStudy (australopithecus):

\[8 \int\limits_{}^{} ((1+\cos(2\theta))/2)^{2}d(\theta)\] trigonometric functions to the even power we use this identity 1 + cos(2theta)/2 Now \[2 \int\limits_{}^{} (1 + 2\cos(2\theta) + \cos^{2}(2\theta))d(\theta)\] Now we can just split the integral \[2( \int\limits_{}^{} 1d(\theta) + \int\limits_{}^{} 2\cos(2\theta)d(\theta) + \int\limits_{}^{} \cos^{2}(2\theta)d(\theta) )\]

OpenStudy (australopithecus):

u = 2(theta) du/2 = d(theta) \[2\int\limits_{}^{}\cos(u)du \] \[2\int\limits_{}^{}\cos^{2}(2\theta)d(\theta) = \int\limits_{}^{}\cos^{2}(u)du = (1/2) \int\limits_{}^{} 1 du + (1/2)\int\limits_{}^{}\cos(2u)du\] = 2theta - 2sin(2theta) + (theta) - sin(4theta)/2 + c

OpenStudy (australopithecus):

and I'm still not done we need to write it in terms of t. I probably made a mistake along the way it is almost certain 2theta - 2sin(2theta) + (theta) - sin(4theta)/2 + c 2t = tan(theta) t = tan(theta)/2 theta = arctan(2t) sec^(2) = (1+4t^2) 2arctan(2t) - 2sin(2arctan(2t)) + arctan(2t) - sin(4arctan(2t))/2 + c = 3arctan(2t) - 2sin(2arctan(2t)) - sin(4arctan(2t))/2 + c is the final answer

OpenStudy (anonymous):

integral of du/(a^2+u^2)=(1/a)arc tan u/a +C let a=1, u=2t so that du=2 dt thus integral of du/(a^2+u^2)=int of 2*2dt/(1+4t^2) =2arc tan 2t +C ......ans